1. The reaction 2A → A2​​​​​ was experimentally determined to be second order with a rate constant, k, equal to 0.0265 M–1min–1. If the initial concentration of A was 4.50 M, what was the concentration of A (in M) after 180.0 min? answer in M.
2. The reaction 2A → A2​​​​​ was experimentally determined to be second order with a rate constant, k, equal to 0.0265 M–1min–1. If the initial concentration of A was 2.50 M, what was the concentration of A (in M) after 180.0 min? answer in M.
3. The reaction 2A → A2​​​​​ was experimentally determined to be second order with a rate constant, k, equal to 0.0265 M–1min–1. If the initial concentration of A was 2.75 M, what was the concentration of A (in M) after 180.0 min? answer in M.
1
Expert's answer
2020-03-27T09:02:05-0400

"-dA/dt=kA^2 \implies _{A_o}^{A_t}\int dA/A^2=_0^t\int(-k)dt"

Solving this, we get;

"1\/A_t-1\/A_0=kt"

This is the second order rate equation.


1.Given : "k=0.0265 M^{\u20131}min^{\u20131}; A_0= 4.50 M"

Thus, "1\/A_t=2\/9+180*0.0265=4.999"

"\implies A_t=0.2M."

2. Given : "k=0.0265 M^{\u20131}min^{\u20131}; A_0= 2.50 M"

Thus, "1\/A_t=2\/5+180*0.0265=5.17"

"\implies A_t=0.1934M."

3.Given : "k=0.0265 M^{\u20131}min^{\u20131}; A_0= 2.75 M"

Thus, "1\/A_t=2\/11+180*0.0265=4.952"

"\implies A_t=0.2019M."